could you please help me solve this Calculate 6 ÷ 1/10

Found 2 solutions by AlgebraLady88, richwmiller:
Answer by AlgebraLady88(44)

6 divide by 1/10
= 6 * 10= 60
Explanation: When you divide by fractions, always find the reciprocal of the second fraction; in this problem the second fraction is 1/10. The reciprocal of that is 10.
Then multiply the first term or fraction( you can alternately write '6' as 6/1) with the reciprocal and you've got your answer.

Answer by richwmiller(17219)

6 ÷ 1/10
One way to think about it is:
we know that 1/10 goes in 1 ten times since 10/10=1
if this is so then 1/10 will go into six 60 times
since we have 10 from 1 ÷ 1/10=10 and 6 so that 6*10=60